# Generative ensemble deep learning severe weather prediction from a deterministic convection-allowing model

### v1

@@ 4.3 Evaluations of uncertainty quantification
The combination of good discard test performance and poor spread-skill reliability performance suggests that all methods have their estimated uncertainties correlated with the prediction errors, but not with the actual spread. That said, the ensemble spread would increase for more difficult prediction cases, but the amount of such increase is not scaled perfectly to the prediction error.
@@ 4.4 Evaluations of CGAN outputs
We have two purposes for conducting this part of the evaluation. First, pattern correlations can show the physics-based consistency of CGAN outputs. For example, if the CGAN generates high CREF values, it is likely that it will generate high 2-m dewpoint values at the same location because strong hydrometer signals typically exist in a moist environment. Second, pattern correlations can reveal the impact of noise within the CGAN-generated samples. Given that the starting point of the sample-generation process is HRRR forecasts randomized with strong Gaussian noise [i.e., N(0, 0.5) compared to the normalized HRRR forecasts, which has standard deviations equal to or lower than 1.0; Section3.3.1], evaluating pattern correlations can make sure that the input noise does not damage the CGAN outputs.
@@ 5 Discussion
This research is the first that applies CNNs and deep generative models to the ensemble post-processing of severe weather. Three research questions were raised in Section 1. The first question related to the performance of CNN-based severe prediction compared to simpler methods. Here, two CNN-based methods were verified and compared to an MLP baseline. Results indicate that the two CNN-based methods produced more skillful predictions with higher BSSs; they also have better resolution by predicting higher probabilities in those correctly verified severe weather cases; this outperformance is especially evident for 00-12Z forecasts. That said, CNN-based severe predictions are at least comparable to other state-of-the-art methods, and they are potentially better for short forecast lead times when the CAM inputs are skillful. An important technical choice that differs from prior works and leads to the success of CNNs is the decoupling of representation learning and classification. Severe weather prediction is a long-tailed classification problem with non-severe weather cases being the head class that accounts for more than 98% of the samples. Without proper treatments, CNNs can overfit to the head class and ignore the tail class. The decoupling of representation learning and classification gives CNNs clearer learning goals in each of their inference stages. The majority of their hidden layers are trained to learn storm-scale representations from the data, whereas the classification layer can handle the sample imbalance by shifting its decision boundaries. We think CNNs, or other deep learning models that can process gridded inputs on an end-to-end basis, can potentially perform well for post-processing CAM output.

### v2

@@ 4.3 Evaluations of uncertainty quantification
The combination of good discard test performance and poor spread-skill reliability performance suggests that all methods can rank the uncertainty of various prediction cases. However, they have underestimated the uncertainty of their predictions. That said, their ensemble spread would increase for more difficult prediction cases, but the amount of such increase is not scaled one-to-one for the prediction error.
@@ 4.4 Evaluations of CGAN outputs
We have two purposes for conducting this part of the evaluation. First, pattern correlations can show the physics-based consistency of CGAN outputs. For example, if the CGAN generates high CREF values, it is likely that it will generate high 2-m dewpoint values at the same location because strong hydrometer signals typically exist in a moist environment. Second, pattern correlations can reveal the impact of noise within the CGAN-generated samples. Given that the starting point of the sample-generation process is HRRR forecasts randomized with strong Gaussian noise [i.e., $\mathcal{N}(0,0.5)$ compared to the normalized HRRR forecasts, which has standard deviations equal to or lower than 1.0; Section3.3.1], evaluating pattern correlations can make sure that the input noise does not damage the CGAN outputs.
@@ 5 Discussion
This research applies CNNs and deep generative models to the ensemble post-processing of severe weather. Three research questions were raised in Section 1. The first question related to the performance of CNN-based severe prediction compared to simpler methods. Here, two CNN-based methods were verified and compared to an MLP baseline. Results indicate that the two CNN-based methods produced more skillful predictions with higher BSSs; they also have better resolution by predicting higher probabilities in those correctly verified severe weather cases; this outperformance is especially evident for 00-12Z forecasts. That said, CNN-based severe predictions are at least comparable to other state-of-the-art methods, and they are potentially better for short forecast lead times when the CAM inputs are skillful. An important technical choice that differs from prior works and leads to the success of CNNs is the decoupling of representation learning and classification. Severe weather prediction is a long-tailed classification problem with non-severe weather cases being the head class that accounts for more than 98% of the samples. Without proper treatments, CNNs can overfit to the head class and ignore the tail class. The decoupling of representation learning and classification gives CNNs clearer learning goals in each of their inference stages. The majority of their hidden layers are trained to learn storm-scale representations from the data, whereas the classification layer can handle the sample imbalance by shifting its decision boundaries. We think CNNs, or other deep learning models that can process gridded inputs on an end-to-end basis, can potentially perform well for post-processing CAM output.
@@
Future work could extend in several directions. First, the CGANs can participate in the training of the CNN-based severe weather prediction model as data augmentation. This idea was not implemented in this research because we intended to measure the contribution of CGANs in the inference stage—when the CGAN ensemble and the CNN baseline were operated on the same CNN model, their performance difference can be attributed to the use of CGAN outputs. Second, CGANs can be improved to generate CAM fields with larger variations. Here, two CGANs were applied to generate 10 predictors by taking 5 explicit storm-scale predictors, including 2-5 km UH, as conditional inputs. This configuration makes it easy for CGANs to generate convective cells in the same locations as their corresponding deterministic CAM forecasts, and it avoids the problem of creating non-existent storms under a stable environment. However, in situations where the CAM forecasts exhibit positional errors, the CGAN outputs may amplify that error. Developing better initial conditions for the sample generation process of CGAN may solve this problem. Finally, the purpose of the CGAN used here was to generate synthetic predictors from a deterministic CAM run. It offers a way to approximate the variability of the forecast data and is not tied to a specific post-processing model or region. For future research, other ML-based severe weather prediction models, e.g., dense neural networks and decision trees, can be applied to form an ensemble learning system. The method that combines CGANs and severe weather post-processing models can also potentially be generalized to other regions where high-quality severe weather observations are available.